I need to use the five-step method to solve this problem. The five steps are:
1. State what you need to solve for.
2. Identify the unknowns.
3. Write the equation derived from the word problem.
4. Solve for the given variable.
5. Check.
This is the question.
One-third of a number is 5 less than half of the same number. What is the number?
(I always get confused on word problems) This is what I think. I don't know for sure.
1. The number must be found.
2. The unknowns are x.
3. 1/3x=5-(1/2)x
4. If this is the correct formula, this is how I solved it.
6((1/3)x + (1/2)x) = (5)6
6(1/3)x + 6(1/2)x = 30
2x + 3x =30
5x =30
x =6
5. Check. (1/3)(6)=5-(1/2)(6)
2=5-3
2=2
Is this the correct formula for this problem? I think I solved it correctly. I just went back to school, so this is actually one of my first word problems I solved correctly, if it is correct?
Thanks alot for checking it. :)

Answer by ewatrrr(24785)

Hi

1. The number must be found.
2. The unknown is x.
3. 1/3x = (1/2)x - 5 | 5 less than half of the same number
4. 2x = 3x - 30
30 = x
5. (1/3)30 = 10 = (1/2)30 - 5
What has been demonstrated here is that 'checking Your answer' is of
no avail, if the question has been algebraically Set-Up incorrectly.
Good work, in general, however. You will be fine, just be very heads up
on the Set Up.